For each quadratic function, state the vertex and then graph the function
The cost equation used by the R & E Manufacturing Co. Is C(h) =2h^2 - 8h + 12, where h is the number of hours it takes to produce a particular product and C(h) Isa the cost of procuction in hundreds of dollars. What is the minimum cost possible?

Answer by stanbon(75887)

For each quadratic function, state the vertex and then graph the function
The cost equation used by the R & E Manufacturing Co. Is C(h) =2h^2 - 8h + 12, where h is the number of hours it takes to produce a particular product and C(h) Is the cost of procuction in hundreds of dollars. What is the minimum cost possible?
-------
Minimum occurs when x = -b/(2a) = 8/(2*2) = 2
-----------------------
C(2) = 2*4-8*2+12 = 4
---------------
Cheers,
Stan H.
---------------